# T-matrix representation of optical scattering response: Suggestion for a data format

### v1

@@ Abstract
keywords:
Data Format , Data Science , Data Storage and Management , Multipolar representations , Scattering , T-matrix
@@ 3 T-matrix formalism
$\langle\sigma_{\text{ext}}\rangle=-\frac{2\pi}{k^{2}}\mathrm{Re}\sum_{l=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{i=1}^{2}T^{ii}_{l,m,l,m}\,,$ (9)
@@
$\langle\sigma_{\text{sca}}\rangle=\frac{2\pi}{k^{2}}\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=1}^{2}\sum_{j=1}^{2}T^{ij}_{l,m,l^{\prime},m^{\prime}}\,,$ (10)
where $k$ is the wavenumber, $l\in\mathbb{N}_{0}$, $m\in\mathbb{Z}$ with $l\geq|m|$, and $i,j$ indices indicate the polarization of the incident and scattered modes. These expressions were derived assuming illumination with plane waves. In the general case, the cross-sections are only proportional to the right-hand side of the equations. The cross-sections are always positive-defined.
@@
.1 hdf5 file. .2 name, description, keywords, storage_format_version. .2 tmatrix. .2 modes. .3 l. .3 m. .3 polarization. .2 frequency. .2 embedding. .3 name, description, keywords . .3 relative_permittivity. .3 relative_permeability. .2 scatterer. .3 material. .4 name, description, keywords . .4 relative_permittivity. .4 relative_permeability. .3 geometry. .4 name, description, keywords, shape. .4 . .2 computation. .3 software, method, name, description, keywords. .3 files. .3 mesh.XYZ. .3 method_parameters.       The required entries given in the following subsections should not be omitted, unless it is mentioned explicitly, since they provide important information concerning the T-matrix and the scatterer of interest.
@@ 2.4 Modes
Finally, for the special case of a cluster of scatterers when the T-matrix is defined in the local basis of each scatterer, the datasets /modes/positions and /modes/index should be added to provide the correspondence between the T-matrix entries and the local coordinate systems.
@@ 2.6.1 Material
The material is a subgroup of the group /NAME describing the scatterer. In case of multiple scatterers, each material subgroup is inside the corresponding scatterer group. The surrounding medium is not included in the scatterer group. The description of the material is typically done by providing the permittivity and permeability or by defining the refractive index and the impedance. Each material group can be annotated with a name, description, and keywords as attribute. The recommended way to add the name attribute of the material is to specify both the chemical formula and the common name if available (e.g., “Au, Gold”). The definitions here assume linear, homogeneous materials. The specific case of a layered structure can be accommodated by entering the material parameters as arrays, and separating the attributes corresponding to each layer with a semicolon in the string. The materials can be dispersive. The keywords can also contain the type of material, e.g., dielectric. Special information, such as nonlocality, can be added as a keyword, and the details can be included in the description. Then, the local permittivity contribution is included into the standard datasets introduced below for material parameters. If the values for the relative permittivity and permeability were used from an external source, this can be specified in the reference attribute, in the form of a reference paper or other link as a string. Alternatively, if the values were measured experimentally, the original datasets should be added under the subgroup /NAME/material/experimental_data, which includes the material parameters as described below and the corresponding frequencies/wavelengths with specified units. The method used for interpolation should be added as well in a string format as an attribute interpolation. We separate cases with isotropic and anisotropic media, as well as cases with and without magnetoelectric coupling. This leads to the four classes of isotropic, biisotropic, anisotropic, and bianisotropic materials, that we describe below (see also A).
@@ 2.6.2 Geometry
If the scatterer has a simple geometric shape, the parameters from Table 2 can be used to describe them. The shape should then be specified by adding the attribute shape. Note that by default for the rotationally symmetric objects, the symmetry axis is the z-axis. In the table, core-shell sphere is separated into a separate entry, but for a general case of a layered scatterer, it is admissible to set the final shape of the scatterer as the shape attribute and add an array of geometrical parameters together with an array of material properties corresponding to each layer. The accepted convention is to measure the radius of the shell from the center of the whole object [82, 83]. If the same unit can be used to describe all geometrical parameters, it can be specified as an attribute unit of the group. The unit can also be an attribute of each individual parameter.
@@
Table 2: Basic three-dimensional shapes and their defining parameters
@@
Cone radius, height
@@
Rectangular cuboid lengthx, lengthy, lengthz
@@
Cut sphere [34] radius, height
Core-shell sphere radius_0 , radius_N
For a general representation of the scatterer, a mesh file has to be defined. The only possible reason for a mesh to be completely absent in the file is the choice of a semi-analytical method for computations (e.g., EBCM, Mie), then the presence of the attribute shape and the corresponding datasets of geometrical parameters are obligatory. The mesh can be specified in /NAME/geometry/mesh.XYZ or /computation/mesh.XYZ, where XYZ stands for the file extension, depending on what is physically more reasonable. For example, for multiple scatterers and a single mesh file, it is sufficient to specify the mesh in /computation. The mesh should be ideally in a common format, e.g., msh or STL. However, STL can only define surface meshes and no volumetric ones, so for complicated structures, such as ones consisting of multiple materials, other formats can be used. To simplify access to the mesh for the user, a softlink /mesh to the mesh location can be provided at the top level, if possible. In exceptional cases, when there is no possibility to describe the mesh using a common format without losing information about the mesh or deforming it, the specific list of mesh parameters datasets can be stored inside a group mesh. A unit of length has to be added to the mesh definition as an attribute unit. The attributes name, description, and keywords can provide additional information.
@@ 4.3 Physics constraints
A rotationally symmetric object produces a scattering response that is also rotationally invariant. This sets constraints on the T-matrix of such an object (Eqs. 30 and 31 in [9]):
@@
$T^{ij}_{l,m,l^{\prime},m^{\prime}}$ $=(-1)^{i+j}T^{ij}_{l,-m,l^{\prime},-m^{\prime}}\,.$ (23)
For a sphere, this condition implies a matrix with only diagonal entries in parity basis.
@@
$T^{ij}_{l,m,l^{\prime},m^{\prime}}=(-1)^{m+m^{\prime}+l+l^{\prime}}\,T^{ij}_{l,m,l^{\prime},m^{\prime}}\,.$ (26)
@@
For a non-absorbing object, one can derive the condition for the corresponding T-matrix by applying the principle of conservation of energy (Eq. 45 in [9]):
@@
If losses are present, but no gain is observed, a more general condition for the T-matrix is fulfilled, i.e., the following expression has to be Hermitian positive-semidefinite (Eq. 10 in [105]):
@@
$\sigma=\frac{1}{2}\frac{\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=1}^{2}\sum_{j=1}^{2}{\lvert T^{ij}_{l,m,l^{\prime},m^{\prime}}-(T^{\prime})^{ij}_{l,m,l^{\prime},m^{\prime}}}\rvert^{2}}{\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=1}^{2}\sum_{j=1}^{2}\lvert{T^{ij}_{l,m,l^{\prime},m^{\prime}}\rvert}^{2}+\lvert{(T^{\prime})^{ij}_{l,m,l^{\prime},m^{\prime}}\rvert}^{2}}\,.$ (29)
@@
11cyl = treams.io.load_hdf5("cylinder_tio2.h5")
@@
24with h5py.File("two_cylinders.h5", "w") as fobj:
@@ Appendix C Mode normalization
Here, we comprehensively define the normalization of the modes starting from elementary functions, which coincides with the normalization in [63]. This is necessary to have an unambiguous definition of the vector spherical waves we use. We also define a reference T-matrix of an object which can be used to verify the normalization. Please note, we work her with complex-valued VSWF. If real-valued expressions are needed, dedicated conversion tools can easily be set up.
@@ Appendix G Full code of the example in TiO2 cylinder
11cyl = treams.io.load_hdf5("cylinder_tio2.h5")
@@
24with h5py.File("two_cylinders.h5", "w") as fobj:
@@ Appendix H Alphabetical list of pre-defined groups, datasets and attributes
/angular_frequency ** array : float, complex
@@
/angular_vacuum_wavenumber ** array : float, complex
@@
/embedding/chirality float, complex
@@
/embedding/non-reciprocity float, complex
@@
/embedding/refractive_index ** array: float, complex; single float, complex
@@
/embedding/relative_impedance ** array: float, complex; single float, complex
/embedding/relative_permeability ** array: float, complex; single float, complex
@@
/embedding/relative_permittivity ** array: float, complex; single float, complex
@@
/mesh or /mesh.XYZ .msh, .STL, etc
@@
/modes/l_incident ** array: int
/modes/m_incident ** array: int
@@
/modes/l_scattered ** array: int
/modes/m_scattered ** array: int
@@
/NAME/geometry/mesh or /NAME/geometry/mesh.XYZ ** (entries containing mesh: same with /computation/ possible ) .msh, .STL, etc.
/NAME/geometry/mesh/file_extension str
/NAME/geometry/mesh/unit * str
/NAME/geometry/mesh.XYZ/unit * .msh, .STL
/NAME/geometry/mesh or /NAME/geometry/mesh.XYZ ** .msh, .STL, etc.
/NAME/geometry/mesh/file_extension str
/NAME/geometry/mesh/unit * str
/NAME/geometry/mesh.XYZ/unit * str
@@
/NAME/material/chirality float, complex
@@
/NAME/material/non-reciprocity float, complex
@@
/NAME/material/refractive_index ** array: float, complex; single float, complex
@@
/NAME/material/relative_impedance ** array: float, complex; single float, complex
@@
/NAME/material/relative_permeability ** array: float, complex; single float, complex
@@
/NAME/material/relative_permittivity ** array: float, complex; single float, complex
@@
/vacuum_wavelength ** array: float, complex; single float, complex
/vacuum_wavelength/unit * array: float, complex; single float, complex
/vacuum_wavenumber ** array: float, complex; single float, complex

### v2

@@ 3 T-matrix formalism
$\langle\sigma_{\text{ext}}\rangle=-\frac{2\pi}{k^{2}}\mathrm{Re}\sum_{l=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{i=\pm 1}T^{ii}_{l,m,l,m}\,,$ (9)
@@
$\langle\sigma_{\text{sca}}\rangle=\frac{2\pi}{k^{2}}\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=\pm 1}\sum_{j=\pm 1}\lvert T^{ij}_{l,m,l^{\prime},m^{\prime}}\rvert^{2}\,,$ (10)
where $k$ is the wavenumber, $l\in\mathbb{N}_{0}$, $m\in\mathbb{Z}$ with $l\geq|m|$, and $i,j$ indices indicate the polarization of the incident and scattered modes, respectively, and take the values of -1 and 1. These expressions were derived assuming illumination with plane waves. In the general case, the cross-sections are only proportional to the right-hand side of the equations. The cross-sections are always positive-defined.
@@
.1 hdf5 file. .2 name, description, keywords, storage_format_version. .2 tmatrix. .2 modes. .3 l. .3 m. .3 polarization. .2 frequency. .3 unit. .2 embedding. .3 name, description, keywords . .3 relative_permittivity. .3 relative_permeability. .2 scatterer. .3 material. .4 name, description, keywords . .4 relative_permittivity. .4 relative_permeability. .3 geometry. .4 name, description, keywords, shape, unit. .4 . .2 computation. .3 software, method, name, description, keywords. .3 files. .3 mesh.XYZ. .3 method_parameters.       The required entries given in the following subsections should not be omitted, unless it is mentioned explicitly, since they provide important information concerning the T-matrix and the scatterer of interest.
@@ 2.4 Modes
Finally, for the special case of a cluster of scatterers when the T-matrix is defined in the local basis of each scatterer, the datasets /modes/positions and /modes/index should be added to provide the correspondence between the T-matrix entries and the local coordinate systems. The index of incident and scattered modes can also be separated into /modes/index_incident and /modes/index_scattered.
@@ 2.6.1 Material
The material is a subgroup of the group /NAME describing the scatterer. In case of multiple scatterers, each material subgroup is inside the corresponding scatterer group. The surrounding medium is not included in the scatterer group. The description of the material is typically done by providing the permittivity and permeability or by defining the refractive index and the impedance. Each material group can be annotated with a name, description, and keywords as attribute. The recommended way to add the name attribute of the material is to specify both the chemical formula and the common name if available (e.g., “Au, Gold”). The definitions here assume linear, homogeneous materials. The specific case of a layered structure can be accommodated by entering the material parameters as arrays, and separating the attributes corresponding to each layer with a semicolon in the string. The materials can be dispersive. In each scatterer group, the wavelength dependence should occur along the first axis of the datasets describing the material parameters. The keywords can also contain the type of material, e.g., dielectric. Special information, such as nonlocality, can be added as a keyword, and the details can be included in the description. Then, the local permittivity contribution is included into the standard datasets introduced below for material parameters. If the values for the relative permittivity and permeability were used from an external source, this can be specified in the reference attribute, in the form of a reference paper or other link as a string. Alternatively, if the values were measured experimentally, the original datasets should be added under the subgroup /NAME/material/experimental_data, which includes the material parameters as described below and the corresponding frequencies/wavelengths with specified units. The method used for interpolation should be added as well in a string format as an attribute interpolation. We separate cases with isotropic and anisotropic media, as well as cases with and without magnetoelectric coupling. This leads to the four classes of isotropic, biisotropic, anisotropic, and bianisotropic materials, that we describe below (see also A).
@@ 2.6.2 Geometry
If the scatterer has a simple geometric shape, the parameters from Table 2 should be used to describe it. The shape should then be specified by adding the attribute shape. Note that by default for the rotationally symmetric objects, the symmetry axis is the z-axis. In the table, core-shell sphere is separated into a separate entry, but for a general case of a layered scatterer, it is admissible to set the final shape of the scatterer as the shape attribute and add an array of geometrical parameters together with an array of material properties corresponding to each layer. The accepted convention is to measure the radius of the shell from the center of the whole object [82, 83]. If the same unit can be used to describe all geometrical parameters, it can be specified as an attribute unit of the group. The unit can also be an attribute of each individual parameter.
Clearly, the provided list of basic shapes cannot cover all scatterers that may be of interest, which limits the search capabilities within the future database. To address this, we strongly emphasize the importance of including the mesh in the data file, as detailed later in the section. The previously mentioned name attribute can be considered as a possible entry to specify names for shapes not included in the basic shapes list. Upon reasonable request, new shapes will be added to the current list, which will be published in the GitHub repository.
@@
Table 2: Basic three-dimensional shapes and their defining parameters.
@@
cut_sphere [34] radius, height
core_shell_sphere radius_0 , radius_N
spheroid radiusxy, radiusz
@@
cone radius_top, radius_bottom, height
ring radius_major, radius_minor, height
@@
rectangular_cuboid lengthx, lengthy, lengthz
@@
pyramid n_edges, radius, height, angle, apex_shift
regular_prism n_edges, radius, height, shift
wedge lengthx, lengthy, lengthz, deltax, deltay
convex_polyhedron points
For a general representation of the scatterer, a mesh file has to be defined. The only possible reason for a mesh to be completely absent in the file is the choice of a semi-analytical method for computations (e.g., EBCM, Mie). In this case, semi-analytical should appear in keywords attribute of /computation, and the presence of the attribute shape and the corresponding datasets of geometrical parameters is obligatory. The mesh can be specified in /NAME/geometry/mesh.XYZ or /computation/mesh.XYZ, where XYZ stands for the file extension, depending on what is physically more reasonable. For example, for multiple scatterers and a single mesh file, it is sufficient to specify the mesh in /computation. The mesh should be ideally in a common format, e.g., msh or STL. However, STL can only define surface meshes and no volumetric ones, so for complicated structures, such as ones consisting of multiple materials, other formats can be used. To simplify access to the mesh for the user, a softlink /mesh to the mesh location can be provided at the top level, if possible. In exceptional cases, when there is no possibility to describe the mesh using a common format without losing information about the mesh or deforming it, the specific list of mesh parameters datasets can be stored inside a group mesh. A unit of length has to be added to the mesh definition as an attribute unit. The attributes name, description, and keywords can provide additional information.
@@ 4.3 Physics constraints
A rotationally symmetric object produces a scattering response that is also rotationally invariant. This sets constraints on the T-matrix of such an object, formulated in the following for the case of the symmetry axis aligned with the z-axis (Eqs. 30 and 31 in [9]):
@@
$T^{ij}_{l,m,l^{\prime},m^{\prime}}$ $=ijT^{ij}_{l,-m,l^{\prime},-m^{\prime}}\,.$ (23)
For a sphere, this condition implies a matrix with only diagonal entries in parity basis, which does not depend on azimuthal indices $m$, $m^{\prime}$. Please note the difference in notation compared to the reference, as the polarization indices $i$, $j$ take the values of -1, 1.
@@
$T^{ij}_{l,m,l^{\prime},m^{\prime}}=(-1)^{m+m^{\prime}+l+l^{\prime}}\,T^{-i-j}_{l,m,l^{\prime},m^{\prime}}\,.$ (26)
@@
For a non-absorbing object, the condition for the corresponding T-matrix can be derived by applying the principle of conservation of energy (Eq. 45 in [9], Eq. 5.59 in [102]). The integral of the Poynting vector $\langle\mathbf{S(r)}\rangle\cdot\mathbf{r}$ over a spherical surface at infinity has to vanish (Eq. 5.55 in  [102]). This leads to the unitarity of the S-matrix: $\mathbf{S}^{\dagger}\mathbf{S}=\mathbb{1}$. The corresponding T-matrix representation is then:
@@
If loss is prevalent in the medium, the integral of the Poynting vector over a spherical surface at infinity is negative, as more energy comes in than goes out. From this inequality, it can be concluded that the following expression is Hermitian positive-semidefinite (Eq. 10 in [105]):
@@
For the gain medium, the sign of the inequality is the opposite, thus the condition for positive-semidefineteness is violated.
@@
$\sigma=\frac{1}{2}\frac{\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=\pm 1}\sum_{j=\pm 1}{\lvert T^{ij}_{l,m,l^{\prime},m^{\prime}}-(T^{\prime})^{ij}_{l,m,l^{\prime},m^{\prime}}}\rvert^{2}}{\sum_{l=1}^{l_{\text{max}}}\sum_{l^{\prime}=1}^{l_{\text{max}}}\sum_{m=-l}^{l}\sum_{m^{\prime}=-l^{\prime}}^{l^{\prime}}\sum_{i=\pm 1}\sum_{j=\pm 1}\lvert{T^{ij}_{l,m,l^{\prime},m^{\prime}}\rvert}^{2}+\lvert{(T^{\prime})^{ij}_{l,m,l^{\prime},m^{\prime}}\rvert}^{2}}\,.$ (29)
@@
11cyl = treams.io.load_hdf5("cylinder_tio2.tmat.h5")
@@
24with h5py.File("two_cylinders.tmat.h5", "w") as fobj:
@@ Appendix C Mode normalization
Here, we comprehensively define the normalization of the modes starting from elementary functions, which coincides with the normalization in [63]. This is necessary to have an unambiguous definition of the vector spherical waves we use. We also define a reference T-matrix of an object which can be used to verify the normalization. Please note, we work here with complex-valued VSWF. If real-valued expressions are needed, dedicated conversion tools can easily be set up.
@@ Appendix G Full code of the example in TiO2 cylinder
11cyl = treams.io.load_hdf5("cylinder_tio2.tmat.h5")
@@
24with h5py.File("two_cylinders.tmat.h5", "w") as fobj:
@@ Appendix H Alphabetical list of pre-defined groups, datasets and attributes
/angular_frequency ** array: float, complex; scalar: float, complex
@@
/angular_vacuum_wavenumber ** array: float, complex; scalar: float, complex
@@
/embedding/chirality array: float, complex; scalar: float, complex
@@
/embedding/non-reciprocity array: float, complex; scalar: float, complex
@@
/embedding/refractive_index ** array: float, complex; scalar: float, complex
@@
/embedding/relative_impedance ** array: float, complex; scalar: float, complex
/embedding/relative_permeability ** array: float, complex; scalar: float, complex
@@
/embedding/relative_permittivity ** array: float, complex; scalar: float, complex
@@
/mesh .msh, .STL, etc
/modes/index ** array: int
/modes/index_incident ** array: int
/modes/index_scattered ** array: int
/modes/l_incident ** array: int
/modes/l_scattered ** array: int
@@
/modes/l_incident ** array: int
/modes/l_scattered ** array: int
@@
/modes/m_incident ** array: int
/modes/m_scattered ** array: int
@@
/modes/positions ** array: float
/NAME/geometry/euler_angles array: float
/NAME/geometry/expansion_center array: float
@@
/NAME/geometry/mesh/ or /NAME/geometry/mesh.XYZ ** (entries containing mesh: same with /computation/ possible ) .msh, .STL, etc.
/NAME/geometry/mesh.XYZ/file_extension str
/NAME/geometry/mesh/unit or /NAME/geometry/mesh.XYZ/unit * str
@@
/NAME/material/chirality array: float, complex; scalar: float, complex
@@
/NAME/material/non-reciprocity array: float, complex; scalar: float, complex
@@
/NAME/material/refractive_index ** array: float, complex; scalar: float, complex
@@
/NAME/material/relative_impedance ** array: float, complex; scalar: float, complex
@@
/NAME/material/relative_permeability ** array: float, complex; scalar: float, complex
@@
/NAME/material/relative_permittivity ** array: float, complex; scalar: float, complex
@@
/vacuum_wavelength ** array: float, complex; scalar: float, complex
/vacuum_wavelength/unit * array: float, complex; scalar: float, complex
/vacuum_wavenumber ** array: float, complex; scalar: float, complex